Susan is 4/5 as old as Ronald. 4 years ago, she was 3/4 as old as Ronald. How old is each?

Answer by richwmiller(17219)

s=4r/5
s-4=3*(r-4)/4
r=20
s=16
check
s/r=4/5
16/20=4/5
ok
16-4=3/4*(20-4)
12=3/4*16
12=12
ok